OpenStudy (anonymous):

Find the derivative of f(x) = 7x + 9 at x = 6.

OpenStudy (anonymous):

f(6) = 7*6 + 9 = 42+9 f(6) = 51

OpenStudy (kirbykirby):

CCan you just use the derivative rule? Or must you do this with the limit definition?

hartnn (hartnn):

using first principles? or can you directly use the derivative formulas?

OpenStudy (anonymous):

did i do it right?

hartnn (hartnn):

lol, hi5 kirb :P :)

OpenStudy (kirbykirby):

Hehe

hartnn (hartnn):

hey barbie, this question belongs to the derivative topic!

OpenStudy (anonymous):

I thought that's how you did it but my choices are: 7 9 6 0

hartnn (hartnn):

nice try though :) but its not correct...

OpenStudy (anonymous):

oh okay sorry i couldn't help :/

hartnn (hartnn):

since you're given choices, lets use the easiest way

hartnn (hartnn):

derivative of x is 1 derivative of constant = 0 this is got from the standard relation, \(\Large d/dx (x^n ) = n x^{n-1}\)

hartnn (hartnn):

so derivative of 7x will be 7(1) = 7 and derivative of 9 is 0 overall, derivative of 7x+9 is simply 7 :)

OpenStudy (anonymous):

hey @hartnn can u please help me?

OpenStudy (anonymous):

ohh okay, do you think you could help e walk through one more?

OpenStudy (anonymous):

@hartnn